i need to simplify the expression -18a^6b^-1/3a^6b
Found 2 solutions by stanbon, richwmiller:
Answer by stanbon(75887)

simplify the expression
[-18a^6b^-1]/[3a^6b]
---
= (-18/3)(a^6/a^6)(b^-1/b)
---
= (-6)(1)(1/b^2)
---
= -6/b^2
=============
Cheers,
Stan H.

Answer by richwmiller(17219)

-18a^6b^(-1)/3a^6b
-(18 a^6 /(b^2 (3 a^6))
b^-1 puts the b in the denominator
The a^6 over a^6 cancel out
We are left with -18/3b^2=-6/b^2